Deployment Team Action items

From GridPP Wiki
Jump to: navigation, search

This page now obsolete and replaced by the Operations_Team_Action_items

This is a Wiki area to track deployment actions

Count list for minute taking

Works left -> right. EVO meeting = 1: F2F meeting = 5

IN: Kashif=29 Alessandra=29 Daniela=28 Duncan=27 Mark=28 Sam=28 Wahid=28 Stuart=27 Jeremy=30 Pete=28 Andrew=28 Derek=28 Brian=30

Minutes can be found at: UKI-ROC agenda

Almost OUT! Jens=21, Mingchao=23 (On loan to EGI team)

OUT! Fraser=8.5 (retired hurt); Stephen=4; Steve=4 (transferred for 1M ukp to CERN!); Olivier=11; Andrew=19 (killed by a falling bash script while re-running YAIM); Greig=20 (an LHCb script moved him to CERN); James=23 (moved on to be sheriff); Dug=24 (infected by evil banker virus); Graeme=27

Action list

Action ID Action description Owner Target date Status Date closed Notes

D-090623-05 Organise reviewers for web pages and wiki sections Jeremy 2009-07-07 Progress 20/07/07 asked in the first instance for volunteers. Remainder will be assigned across team to balance. Revisit in November. It is mainly the sysadmins introductory pages that need attention. Update December 2010: This is still ongoing and needs to be brought in line with timelines for website changes for GridPP4. Review at meeting on 8th February 2011.

D-100427-01 Migration strategy for ROC? Jeremy Open 2010-11-02 Current situation appears to be fine; further changes depend on the NGS side of things. December 2010: JG still suggests early in 2011. Late January Ireland have started the validation process, so expect NGI_UK process to start in mid-February.

D-101123-01 Provide input on hardware scaling issues for disk servers at Tier-2s, particularly TB_disk/Gb_networking ratios. Sam, Brian, Wahid 2010-12-07 Open 2011-01-12 Being discussed internally to provide a consensus value.
D-110208-01 Review status of VO share publishing at sites in GStat2. All 2011-02-08 Open

D-110222-04 Review the need for GridPP VO lists Jeremy 2011-03-01 Open 2011-02-22: Revised text which was out of date. PG updated dteam entry. Review of VO information and check on CIC generation of vomses entries ongoing.
D-110412-01 Raise issues with ATLAS Panda service view of experiment software being site, not cluster or queue, resolution. (Which causes huge problems with heterogeneous clustered sites). Alessandra 2011-04-12 Open
D-110503-01 Confirm if LHCb jobs can be restricted to 24hrs at T2s Raja 2011-05-10 Open

See also: Deployment Team Completed Actions See also: Deployment Issues